OpenStudy (anonymous):

In order to solve the following system of equations by subtraction, which of the following could you do before subtracting the equations so that one variable will be eliminated when you subtract them? 4x-2y=7 3x-3y=15 A. Multiply the top equation by 3 and the bottom equation by 4. B. Multiply the top equation by 15 and the bottom equation by 7. C. Multiply the top equation by 1/3. D. Multiply the top equation by -3 and the bottom equation by 2.

OpenStudy (anonymous):

D

OpenStudy (aravindg):

The idea is to make coefficient of x or y's same. Think of the option that helps you do that.

OpenStudy (anonymous):

how do i make coefficients the same? @AravindG

OpenStudy (aravindg):

For instance consider the first option Multiply the top equation by 3 : (4x-2y=7)3=12x-6y=21 Multiply bottom equation by 4. (3x-3y=15)4=12x-12y=60 Notice that we got coefficients of x as same. So when we subtract them x terms get eliminated.